Mike took clothes to the cleaners three times last month.​ First,he brought 4 shirts and 2 pairs of slacks and paid ​$17.94. Then he brought 7 ​shirts, 2 pairs of​ slacks, and 2 sports coats and paid ​$40.39. ​Finally, he brought 3 shirts and 1 sports coat and paid ​$14.96. How much was he charged for each​ shirt, each pair of​ slacks, and each sports​ coat?
Answer by Boreal(15235)

4SH+2S=17.94
7SH+2S+2SC=40.39
3SH+1SC=14.96
---------------------
7SH+2S+2SC=40.39
-4SH-2S=-17.94, multiplying by -1 to eliminate S.
3SH+2SC=22.45
-3SH-1SC=-14.96, multiplying the third by -1 to eliminate SH
1SC=$7.49
Therefore, 3SH=14.96-7.49=7.47
1SH=$2.49
4(2.49)+2S=17.94; 9.96+2S=17.94; 2S=7.98
1S=$3.99
--
shirts are $2.49 each
slacks are $3.99 each
sport coats are $7.49 each
--
Check with the second
7(2.49)+2(3.99)+2(7.49)=17.43+7.98+14.98=$40.39 checks